# TAME: Temporal Audio-based Mamba for Enhanced Drone Trajectory Estimation and Classification

### v4

@@ Abstract
Index Terms— Anti-UAV, Audio, Trajectory Estimation, Classification, Mamba.
1 Introduction
Compact unmanned aerial vehicles (UAVs) [1] are popular for their portability, ease of use, and affordability, making them valuable in fields such as transportation [2, 3], photography [4, 5, 6, 7], search and rescue [8, 9, 10]. However, these same features also make them difficult to detect when used maliciously, posing threats to air traffic control and security and being exploited in warfare and border drug smuggling [11], as shown in Fig. 1.
Existing anti-UAV detection methods often rely on single-modality techniques [12, 13, 14, 15, 16, 17, 18], such as radio signals, radar, or images [19], but these approaches can be error-prone, particularly due to limitations in radar cross section [20, 21, 22] or pixel size [23]. While some methods attempt to overcome these issues through multi-modal fusion [24, 25, 26], combining inputs like multi-spectral images and audio data, they often rely on unrealistic assumptions [27], such as perfect targeting or high vantage points, making them less practical in real-world scenarios. Audio, however, presents a distinct advantage by providing reliable information about the direction [28, 29], classification [30, 31], and distance of sound sources [32, 33], which is largely unaffected by environmental conditions. This is particularly useful given the significant noise generated by UAVs, making them easier to detect through audio [24]. Recognizing these benefits, recent studies [25] have begun to advocate for integrating audio-based solutions to improve UAV detection accuracy.
@@ 2 Method
2.1 Temporal Spectral Mamba
Audio features are extracted by TSMamba, as shown in Fig. 2. Let denote the temporal (t) or spectral (s) axis. Denote as the number of microphone channels, R as the temporal width of the spectrogram, S as the spectral height, W is the width of the patch, and H as the height of the patch. To extract audio features, multi-channel audio is converted into mel-spectrograms , which serve as the input to our model. As the standard Mamba is a 1D sequence model, the spectrogram is split and flattened along the temporal and spectral axis to obtain , which consists of the temporal patch sequence and the spectral patch sequence . Inspired by ViT [35], we append learnable tokens in both temporal and spectral directions. These tokens are utilized by selective SSM to summarize the temporal difference and spectral attenuation features of the entire patch sequence. To utilize those tokens, is linearly projected to a vector of size D, and position embeddings are added. With the linearly projected vector, we can generalize the overall feature vector :
@@ 2.1 Temporal Spectral Mamba
2.2 Temporal Feature Enhancement
@@ 2.3 Detection Head
TAME(Ours) | 6.61 | 0.11 | 0.30 | 0.34 | 0.55 | 98.0 | 0.11 | 0.30 | 0.34 | 0.55 | 98.0 | 0.55 | 98.0 |
Best results are highlighted in bold, and second best in . This notation is used throughout the paper. : the mean value of day and night.
@@ 3.3 Baseline Selections
3.3.1 Trajectory Estimation and UAV Classification
Table 1 shows that TAME achieves state-of-the-art results in both and , outperforming audio, visual, and audio-visual fusion methods, especially at night. TAME’s ability to capture global temporal features in audio spectrograms gives it an edge over other models like DroneChase and AudioNet. While visual-based models excel during the day, their performance declines at night, and audio-visual fusion methods struggle with lighting changes. Although attention mechanisms in AV-PED and AV-FDTI improve performance, they still fall short of TAME. This highlights the superiority of single-modal approaches like TAME in UAV trajectory estimation and classification. Fig. 4 and 5 illustrate UAV position estimates and classification results, showing that most UAVs are accurately detected, except for some cases like M300, where weak audio signals are masked by environmental noise.
@@ 3.3.1 Trajectory Estimation and UAV Classification
0.59 | 97.7 | |||||
0.87 | 95.4 | |||||
0.68 | 96.7 | |||||
0.55 | 98.0 |
3.4 Ablation Study and Analysis
To verify the effectiveness of the proposed TAME, we conducted ablation experiments to assess the significance of various features and fusion methods. Table 2 demonstrates that TAME achieves the highest overall performance in both the and . This improvement is attributed to the use of TFE, which enhances temporal features.
4 Conclusion

### v7

@@ Abstract
Index Terms:
Anti-UAV, Audio, Trajectory Estimation, Classification, Mamba.I Introduction
Compact unmanned aerial vehicles (UAVs) [1] are popular for their portability, ease of use, and affordability, making them valuable in fields such as transportation [2, 3], photography [4, 5, 6, 7], search and rescue [8, 9, 10, 11]. However, these same features also make them difficult to detect [12] when used maliciously, posing threats to air traffic control and security and being exploited in warfare and border drug smuggling [13], as shown in Fig. 1.
Existing anti-UAV detection methods often rely on single-modality techniques [14, 15, 16, 17, 18, 19, 20], such as radio signals [21, 22], radar, or images [23], but these approaches can be error-prone, particularly due to limitations in radar cross section [24, 25, 26] or pixel size [27]. While some methods attempt to overcome these issues through multi-modal fusion [28, 29, 30, 31], combining inputs like multi-spectral images and audio data, they often rely on unrealistic assumptions [32], such as perfect targeting or high vantage points, making them less practical in real-world scenarios. Audio, however, presents a distinct advantage by providing reliable information about the direction [33, 34], classification [35, 36], and distance of sound sources [37, 38], which is largely unaffected by environmental conditions. This is particularly useful given the significant noise generated by UAVs, making them easier to detect through audio [28]. Recognizing these benefits, recent studies [29] have begun to advocate for integrating audio-based solutions to improve UAV detection accuracy.
@@ II Method
II-A Temporal Spectral Mamba
Audio features are extracted by TSMamba, as shown in Fig. 2. Let denote the temporal (t) or spectral (s) axis. Denote as the number of microphone channels, R as the temporal width of the spectrogram, S as the spectral height, W is the width of the patch, and H as the height of the patch. To extract audio features, multichannel audio is converted into mel-spectrograms , which serve as input to our model. As the standard Mamba is a 1D sequence model, the spectrogram is split and flattened along the temporal and spectral axis to obtain , which consists of the temporal patch sequence and the spectral patch sequence . Inspired by ViT [40], we append learnable tokens in both temporal and spectral directions. These tokens are used by selective SSM to summarize the temporal difference and spectral attenuation features of the entire patch sequence. To use those tokens, is linearly projected to a vector of size D with added position embeddings . With the linearly projected vector, we can generalize the overall feature vector :
@@ II-A Temporal Spectral Mamba
II-B Temporal Feature Enhancement
@@ II-C Detection Head
| TAME(Ours) | 6.61 | 0.11 | 0.30 | 0.34 | 0.55 | 98.0 | 0.11 | 0.30 | 0.34 | 0.55 | 98.0 | 0.55 | 98.0 | |
Best results are highlighted in bold, and second best in . This notation apply for rest of the paper. : the mean of day and night.
@@ III-C Baseline Selections
III-C1 Trajectory Estimation and UAV Classification
Table I shows that TAME achieves state-of-the-art results in both and , outperforming audio, visual, and audio-visual fusion methods, especially at night. TAME’s ability to capture global temporal features in audio spectrograms gives it an edge over other models like DroneChase and AudioNet. While visual-based models excel during the day, their performance declines at night, and audio-visual fusion methods struggle with lighting changes. Although attention mechanisms in AV-PED and AV-FDTI improve performance, they still fall short of TAME. This highlights the superiority of single-modal approaches like TAME in UAV trajectory estimation and classification. Fig. 4 and 5 illustrate UAV position estimates and classification results, showing that most UAVs are accurately detected, except for some cases like M300, where weak audio signals are masked by environmental noise.
@@ III-C1 Trajectory Estimation and UAV Classification
| 0.59 | 97.7 | |||||
| 0.87 | 95.4 | |||||
| 0.68 | 96.7 | |||||
| 0.55 | 98.0 |
III-D Ablation Study and Analysis
To verify the effectiveness of the proposed TAME, we conducted ablation experiments to assess the significance of various features and fusion methods. Table II demonstrates that TAME achieves the highest overall performance in both the and . This improvement is attributed to the use of TFE, which enhances temporal features.
IV Conclusion